OpenStudy (anonymous):

2(5x+4)=10x+6 is this equation sometimes always or never true?

OpenStudy (adajiamcneal):

Simplifying 2(5x + -4) = 10x + 6 Reorder the terms: 2(-4 + 5x) = 10x + 6 (-4 * 2 + 5x * 2) = 10x + 6 (-8 + 10x) = 10x + 6 Reorder the terms: -8 + 10x = 6 + 10x Add '-10x' to each side of the equation. -8 + 10x + -10x = 6 + 10x + -10x Combine like terms: 10x + -10x = 0 -8 + 0 = 6 + 10x + -10x -8 = 6 + 10x + -10x Combine like terms: 10x + -10x = 0 -8 = 6 + 0 -8 = 6 Solving -8 = 6

OpenStudy (adajiamcneal):

This equation is invalid, the left and right sides are not equal, therefore there is no solution.